# ConfQA: Answer Only If You Are Confident

### v1

@@ Abstract
Can we teach Large Language Models (LLMs) to refrain from hallucinating factual statements? In this paper we present a fine-tuning strategy that we call ConfQA, which can reduce hallucination rate from 20-40% to under 5% across multiple factuality benchmarks. The core idea is simple: when the LLM answers a question correctly, it is trained to continue with the answer; otherwise, it is trained to admit "I am unsure". But there are two key factors that make the training highly effective. First, we introduce a dampening prompt—"answer only if you are confident"—to explicitly guide the behavior, without which hallucination remains high as 15%-25%. Second, we leverage simple factual statements, specifically attribute values from knowledge graphs, to help LLMs calibrate the confidence, resulting in robust generalization across domains and question types. Building on this insight, we propose the Dual Neural Knowledge framework, which seamlessly select between internally parameterized neural knowledge and externally recorded symbolic knowledge based on ConfQA’s confidence. The framework enables potential accuracy gains to beyond 95%, while reducing unnecessary external retrievals by over 30%.
@@ 1 Introduction
Despite the remarkable capabilities that Large Language Models (LLMs) have demonstrated, hallucination of factual statements remains a challenge (Maynez et al., 2020; Zhou et al., 2021; Ji et al., 2023a). It has been broadly realized that factual information shall not be fabricated or generalized, instead shall be anchored in internally parameterized neural knowledge or externally recorded symbolic content (stored in knowledge graphs, webpages, or other repositories). Significant progress has been made in both knowledge internalization through pre-training (Grattafiori et al., 2024) and external knowledge utilization via Retrieval-Augmented Generation (RAG) (Wei et al., 2021; Yu et al., 2022; Gao et al., 2024; Fan et al., 2024; Huang and Huang, 2024). However, a critical question remains: when should LLMs rely on parameterized knowledge versus external sources? In the absence of external data, we wish the LLM to state only reliably parameterized knowledge and eliminate hallucinations; in the presence of external data, we wish the LLM to conduct retrieval wisely to reduce latency, save resources, and avoid distractions (Jiang et al., 2023).
In this paper, we study this problem through three questions:
  • Q1. Does an LLM know what it knows?
  • Q2. Can we teach LLMs to refrain from hallucinations?
  • Q3. What is the optimal strategy for RAG triggering?
We conducted extensive experimental studies to answer these questions and made three contributions.
Figure 1: Overall factuality improvement of our Dual Neural Knowledge framework. Our fine-tuned model ConfQA reduces hallucination to under 5%; when combined with RAG, we increase accuracy by 45% on average while cutting latency by selective RAG triggering.
@@
Our second contribution is an effective fine-tuning method for hallucination suppression, called ConfQA (Q2). Our training data checks the LLM’s inherent answer to a question, and teaches it to state “I am unsure” if the answer is incorrect. Though seemingly simple, there are two key ingredients to make the training effective. First, we discover an important dampener prompt—“Answer only if you are confident”, which plays a crucial role in guiding the LLM behavior. Second, the training data comprise exclusively of simple questions that inquire about an attribute of an entity; by concentrating the regularization on simple factual statements, which serves as building blocks to more complicated factual statements, the training teaches LLMs to generalize this behavior. ConfQA is able to reduce hallucinations to under 5% on various benchmarks (see Figure 1).
Our method exhibits strong transferability: although trained solely on questions derived from DBPedia, similar improvements are observed across other domains (e.g., IMDb), benchmarks (CRAG (Yang et al., 2024), SimpleQA (Wei et al., 2024a)), and long-form answers (LongFact (Wei et al., 2024b), Biography (Min et al., 2023)). Additionally, the fine-tuned model maintains its performance on general benchmarks such as MMLU (Hendrycks et al., 2021)(Section 5).
Our final contribution is an information-providing framework that we call the Dual Neural Knowledge (DualKnowl) framework (Q3). The framework simultaneously invokes both ConfQA and a RAG pipeline in parallel, but it halts the RAG pipeline early unless one of the two conditions is met: (1) when a question requests dynamic information—facts that may change over minutes, days, or years (Vu et al., 2023), or (2) when the ConfQA model responses with “I am unsure” to a factual question. Our empirical study shows that DualKnowl achieves QA accuracy comparable to always invoking RAG, while reducing latency by over 600ms on CRAG (Yang et al., 2024) (Section 6).
2 Related Work
Hallucination, with a systematic taxonomy introduced in Tonmoy et al. (2024), has been listed as a top open problem for Generative AI (Cao et al., 2023). There are two bodies of work on hallucination reduction: LLM post-training and RAG. We next discuss each in detail.
RAG with external knowledge:
Retrieval-augmented generation (RAG) (Lewis et al., 2020; Shi et al., 2024) has been extensively researched in academia and widely applied in industry to enhance LLMs’ capability to answer factual questions accurately; there have been numerous surveys on RAG, such as  Wei et al. (2021)Yu et al. (2022)Gao et al. (2024)Fan et al. (2024)Huang and Huang (2024). What is closely related to our work is RAG triggering strategies. Su et al. (2024) propose a strategy to trigger RAG when the generated token has high entropy or high self-reported attention. Similarly, Jiang et al. (2023) propose to use LLM internalized knowledge for generation, and apply RAG when the confidence of the next token is low.  Peng et al. (2023) proposes a system that conducts retrieval and answer generation iteratively, revising the LLM prompts to improve model responses using the factuality score from an automatic verifier. These methods all use token-level confidence, and often need multiple retrievals sequentially. Our work focuses on fact-level confidence, can apply in situations when token-level confidence is unavailable, and typically require a single retrieval.
LLM fine-tuning:
Training-based hallucination mitigation has been surveyed in Tonmoy et al. (2024). There are two directions for training: enriching the parameterized knowledge, and suppressing hallucinations.
A lot of study focuses on knowledge enrichment during post-training, such as using ground truths in QA data sets (Zhang et al., 2024), leveraging trustworthy sources and especially knowledge graphs (Bayat et al., 2023; Ji et al., 2023b), constructing factuality preference pairs for DPO (Rafailov et al., 2024) based on long form generation, and distilling from larger models (Elaraby et al., 2023). Recently it was discovered that feeding knowledge in post training would encourage the LLM to ignore what it learns during pre-training, thus leading to even more hallucinations (Lin et al., 2024; Gekhman et al., 2024). This is confirmed in our experiments.
On the other hand, a lot of research focuses fine-tuning on hallucination suppression. Sun et al. (2023b) teaches LLMs to recite factual passages to avoid hallucination. Dhuliawala et al. (2023) verifies responses with internalized knowledge before final generation. Tian et al. (2023) generates factuality preference rankings to favor factual statements consistent with external sources or internal knowledge. Xie et al. (2025) trains a factuality evaluator to provide LLM generators with claim-level factuality feedback. A few works focused on teaching LLMs about its confidence on factual statements, including pre-LLM work (Mielke et al., 2022; Kadavath et al., 2022) and recent work based on local intrinsic dimension (Yin et al., 2024).
The works that are closest to ours are Cheng et al. (2024) and Zhang et al. (2024), both instructing the LLM to refuse to answer questions where it has a low confidence about the answer. R-Tuning (Zhang et al., 2024) identifies such questions according to answer correctness and pads sure or unsure to the end of the answer. Cheng et al. (2024) in addition requires providing the correct answer consistently. Our ConfQA training is tremendously different in two ways: first, we use the dampener prompt, which reduces hallucination further by 5-11% in our empirical study; second, we focus on simple factual questions from the DBPedia knowledge graph, which increases factuality by up to 30%. Our experiments also show that requiring consistency in addition can cause a large correctness regression. We present the experimental comparison in detail in Section 5.
@@ 3 Experiment Setup
As we focus on factuality, we wish to reduce hallucination on factual statements, without regressing performance on general knowledge and problem-solving tasks. We thus consider benchmarks in three categories: 1) Short-form factuality benchmarks, where the answers are mostly short; the question include both simple questions regarding an attribute of an entity, and complex ones that require comparison, aggregation, reasoning, and post-processing; 2) Long-form factuality benchmarks, where answers are expected to be long and contain multiple factual statements; 3) General knowledge benchmarks, which focuses on general knowledge and reasoning in diverse disciplines. Under these three categories we considered seven benchmarks, where Table 1 summarizes the benchmarks and Appendix 10 gives details.
@@
We use slightly different metrics to measure the factuality of answers generated by LLMs. For short-form factuality benchmarks, we compute the percentage of correct, incorrect, and missing (not attempted) answers, and then define two top-line metrics.
  • Factuality-score: Following the CRAG benchmark (Yang et al., 2024), we compute Factuality = correct% - incorrect%, which ranges from -1 to 1.
  • F1-score: Following the SimpleQA benchmark (Wei et al., 2024a), we compute F1-score as a harmonic mean of precision and recall, where Precision = correct% / (correct% + incorrect%), and Recall = correct%. F1-score ranges from 0 to 1.
Between the two metrics, F1-score is more lenient for incorrect answers (hallucinations), but factuality strongly prefers missing answers to hallucinations (see examples in Section 10). We decide the correctness of an answer by prompting Llama-3.1-405B to compare a generated answer with the ground truth. We use SimpleQA evaluation prompt 111 https://github.com/openai/simple-evals for the SimpleQA benchmark, and use Prompt 2 in Appendix 9 for other benchmarks. Similar to Yang et al. (2024), we observe a  99% accuracy in the judgment.
For long-form responses we use the automatic evaluation metric, VeriScore (Song et al., 2024), which computes precision, recall, and F1-score. The main difference from short-form questions is that we set the minimum number of facts required for perfect recall based on the median number of extracted claims per dataset, using their fine-tuned models for claim extraction and verification.
For general knowledge benchmarks, MMLU provides ground truths for the multi-choice questions. The score is computed as the percentage of correctly answered questions, as a weighted average among the 57 diverse subjects.
@@
We conduct experiments using six well-known LLMs: Llama3.1-8B, Llama3.1-70B (Touvron et al., 2023), GPT-4o-mini and GPT-4o (OpenAI, 2023), Claude3.5-Sonnet [^2] https://www.anthropic.com/news/claude-3-5-sonnet and Claude3.5-Haiku [^3] https://www.anthropic.com/claude/haiku. We conducted experiments on Nvidia H100 96GB HBM2e GPUs with different configurations based on different model sizes. For Llama3.1-70B models, we did fine tuning on 32 H100 GPUs and inference on 8 H100 GPUs.
4 Q1. Does an LLM Know What It Knows?
We first investigate whether an LLM can accurately estimate its own confidence in a factual statement. If so, we can rely on this estimation to choose between internalized neural knowledge and external symbolic knowledge. We measure confidence in two ways: self-reported confidence, and consistency of answers. We focus on the three short-form benchmarks: Head-to-Tail, SimpleQA, and CRAG. We investigate the calibration between confidence and QA accuracy; that is, when the model has a confidence of 0.8, is the QA accuracy close to 80%?
Figure 2: LLMs’ self-reported confidences is correlated with QA accuracy, but often over-confident. The answer consistency is often better calibrated with QA accuracy.
Confidence vs. Accuracy:
To obtain self-reported confidence, we prompt the LLM to directly provide a confidence score between 0 and 1 along with its answer (prompt template in Appendix 3). We remove missing answers, divide the reported confidences by equal-sized quantile bins, and plot the average accuracy within each bin. The top panels in Figure 2 show the calibration, leading to four observations. (We in addition compare calibration for head, torso, tail entities in Appendix 11).
@@
Consistency vs. Accuracy:
To measure consistency, we ask LLM the same question 20 times with the temperature set to 1.0, select the most frequent response as the final answer, and calculate its frequency among the 20 times as the consistency score. To be robust against minor differences, we determine the "most frequent" answer based on semantic similarity rather than exact string match.
The bottom panels in Figure 2 shows that consistency is mostly better calibrated than self-reported confidence, and largely overlays with the ideal calibration on CRAG. In addition, the calibration curve is more linear compared to self-reported confidence.
Summary:
Ideally, an LLM would know what it knows, give facts it has high confidence about, and refrain from including low-confidence facts to avoid hallucinations. Unfortunately, as this analysis shows, self-reported confidence tends to be over-confident, whereas answer consistency requires invoking LLMs multiple times and thus is impractical. We thus need to find a different strategy.
5 Q2. Can We Teach LLMs to Refrain from Hallucinating?
In this section, we shift gears and study whether we can fine tune an LLM to suppress hallucinations, by only answering a question that it has high confidence about, and admitting “I am unsure” otherwise. We have three key intuitions for this fine-tuning. First, the calibration study shows that LLM’s self-reported confidence is associated with answer accuracy but not calibrated, we thus teach it to calibrate its confidence by showing the ground truth. Second, we introduce a dampener prompt “Answer only if you are confident”, to explicitly guide LLM’s behavior. Third, as we wish to regularize behavior only for factual statements, we focus the teaching on atomic facts (attributes of entities) to avoid distractions of other factors.
5.1 Fine-tuning setup
Training:
We conducted supervised fine-tuning (SFT) and constructed the training data as a set of question-label pairs, where the label teaches the model the answer it shall generate.
The questions ask for atomic facts, and are generated from DBPedia, which covers a diverse set of domains (Intuition #3). We used the open-sourced script from Sun et al. (2023a) to generate question-answer pairs from DBPedia (exclusive from those in the Head-to-Tail benchmark), evenly distributed across head, torso, and tail entities.
We generated labels as follows. First, we prompted Llama-3.1-70B model to answer the DBPedia-based questions (Prompt 1 in Appendix 9). Then, we prompt Llama-3.1-405B to judge if the answer is consistent with the ground truth (Prompt 2 in Appendix 9). If the answer is correct, the label is the ground truth answer; otherwise, the label is “I am unsure about the answer” (Intuition #1).
Inferences:
We call our fine-tuned model ConfQA. At inference time, we prompt QA in two ways: without the dampener, and with the dampener, to examine the dampening effect. (Intuition #2).
Other solutions:

### v2

# ConfRAG: Confidence-Guided Retrieval-Augmenting Generation

@@ Abstract
Can Large Language Models (LLMs) be trained to avoid hallucinating factual statements, and can Retrieval-Augmented Generation (RAG) be triggered only when necessary to reduce retrieval and computation costs? In this work, we address both challenges simultaneously. We introduce ConfQA, a fine-tuning strategy that reduces hallucination rates from 20–40% to below 5% across multiple factuality benchmarks. The approach is simple: when the model answers correctly, it is trained to output the answer; otherwise, it is trained to respond with “I am unsure.” Two design choices make this training effective: (1) a dampening prompt (“answer only if you are confident”) that explicitly discourages overconfident hallucinations, and (2) training data drawn from atomic factual statements (e.g., knowledge graph attribute values), which calibrates model confidence and yields robust generalization across domains and question types. Building on ConfQA, we propose ConfRAG, a triggering strategy that invokes RAG only when the model responses with unsure. This framework achieves accuracy above 95% in ideal case while reducing unnecessary external retrievals by over 30%.
@@ 1 Introduction
Despite the remarkable capabilities that Large Language Models (LLMs) have demonstrated, hallucination of factual statements remains a challenge (Maynez et al., 2020; Zhou et al., 2021; Ji et al., 2023). It has been broadly realized that factual information shall not be fabricated or generalized, instead shall be anchored in internally parameterized neural knowledge or externally recorded symbolic content (stored in knowledge graphs, webpages, or other repositories). Significant progress has been made in both knowledge internalization through pre-training (Grattafiori et al., 2024) and external knowledge utilization via Retrieval-Augmented Generation (RAG) (Wei et al., 2021; Yu et al., 2022; Gao et al., 2024; Fan et al., 2024; Huang and Huang, 2024). However, a critical question remains: when should LLMs rely on parameterized knowledge versus external sources?
Existing RAG-triggering strategies tend to fall at two extremes. In academic settings, triggering is often decided at the token level, for example when a generated token exhibits high entropy, high self-reported attention, or low confidence (Su et al., 2024; Jiang et al., 2023). This approach requires close monitoring of hidden-state signals derived from LLM’s intermediate activations. In contrast, industry practice typically relies on manual or coarse-grained triggers—for instance, enabling RAG only when a user selects a particular model version, when the query falls into certain domains, or when the requested information is known to change over time. However, as highlighted in Head-to-Tail (Sun et al., 2024), LLMs can still make mistakes even when answering static questions, especially for entities of torso-to-tail popularity, and such errors happen across domains.
In this paper, we focus on the problem of determining when to trigger RAG for questions that seek static information. The underlying intuition is straightforward: if an LLM can accurately assess its own knowledge, it should only consult external sources when it recognizes uncertainty. However, our experiments on three benchmarks reveal that self-reported confidence is systematically overestimated (confirming observation in Wei et al. (2024a)) and therefore unreliable for RAG-triggering decisions (Figure 3). To address this, we propose to explicitly teach LLMs to estimate their confidence in factual responses and to use this calibrated confidence as the basis for triggering RAG. In particular, we make the following three contributions.
Figure 1: Factuality improvement of ConfRAG and fine-tuned model ConfQA.
Figure 2: ConfRAG invocation architecture.
Our first contribution is a RAG triggering strategy called ConfRAG, which formulates the triggering decision as detecting low-confidence answers. See Figure 2 for the architecture. For static factual questions, the framework runs LLM-based QA and a RAG pipeline in parallel, but halts the RAG process early if the LLM responses with “I am unsure”. Unlike approaches that rely on inspecting hidden-state signals, this strategy is lightweight and broadly applicable.
Our second contribution is an effective fine-tuning method for confidence recognition, called ConfQA, which forms the core of our triggering strategy. Our training procedure checks the LLM’s inherent answer to a question, and teaches it to state “I am unsure” when the answer is incorrect. Though seemingly simple, two key design choices make it effective. First, we introduce a dampening prompt—“Answer only if you are confident”, which plays a crucial role in shaping the LLM’s behavior. Second, the training data comprise exclusively of simple factual questions about entity attributes; by focusing the regularization on such atomic facts, which serve as building blocks for more complicated statements, the training enables LLMs to generalize this confidence behavior to broader domains.
Our final contribution is a comprehensive empirical study across 7 benchmarks spanning short-form QA, long-form QA, and general knowledge QA. We show that, in theory, ConfRAG can achieve over 95% accuracy with a perfect RAG system; with a real RAG system, ConfRAG attains QA accuracy comparable to always invoking RAG, while reducing P50 latency by over 600ms on CRAG (Yang et al., 2024). Our experiments also reveal two standalone uses of ConfQA when RAG is not yet an option (such as in Speech-in Speech-out systems (Xie and Wu, 2024; Nguyen et al., 2022)): for maintaining correctness with reduced hallucinations, we recommend ConfQA without the dampener in inference, which preserves accuracy with a mild reduction in hallucinations; for minimizing hallucinations, we recommend ConfQA with the dampener, which reduces hallucinations to under 5% (see Figure 2).
2 Related Work
There are two bodies of work related to our work: RAG triggering and LLM post-training. We next discuss each in detail.
RAG with external knowledge:
RAG has been extensively researched in academia and widely applied in industry to enhance LLMs’ capability to answer factual questions accurately; there have been numerous surveys on RAG, such as Wei et al. (2021), Yu et al. (2022), Gao et al. (2024), Fan et al. (2024), Huang and Huang (2024). What is closely related to our work is RAG triggering strategies. Su et al. (2024) propose a strategy to trigger RAG when the generated token has high entropy or high self-reported attention. Similarly, Jiang et al. (2023) propose to use LLM internalized knowledge for generation, and apply RAG when the confidence of the next token is low. Peng et al. (2023) proposes a system that conducts retrieval and answer generation iteratively, revising the LLM prompts to improve model responses using the factuality score from an automatic verifier. Self-RAG (Asai et al., 2023) trains a model to adaptively retrieves documents and do self-reflection using special tokens. These methods all use token-level confidence, and often need multiple retrievals sequentially. Our work focuses on fact-level confidence, can apply in situations when token-level confidence is unavailable, and typically require a single retrieval.
LLM hallucination suppression:
Training-based hallucination mitigation has been surveyed in Tonmoy et al. (2024). There are two directions for training: enriching the parameterized knowledge, and suppressing hallucinations, the latter more related to our work. Sun et al. (2023b) teaches LLMs to recite factual passages to avoid hallucination. Dhuliawala et al. (2023) verifies responses with internalized knowledge before final generation. Tian et al. (2023) generates factuality preference rankings to favor factual statements consistent with external sources or internal knowledge. Xie et al. (2025) trains a factuality evaluator to provide LLM generators with claim-level factuality feedback. Grattafiori et al. (2024) incorporates refusals in training data for samples that got consistently informative and incorrect responses, similar as in Cheng et al. (2024). A few works focused on teaching LLMs about its confidence on factual statements, including pre-LLM work (Mielke et al., 2022; Kadavath et al., 2022; Cohen et al., 2024) and recent work based on local intrinsic dimension (Yin et al., 2024).
The works that are closest to ours are Cheng et al. (2024) and Zhang et al. (2024a) (also very similar to idea in Kapoor et al. (2025)), both instructing the LLM to refuse to answer questions where it has a low confidence about the answer. R-Tuning (Zhang et al., 2024a) identifies such questions according to answer correctness and pads sure or unsure to the end of the answer. Cheng et al. (2024) in addition requires providing the correct answer consistently. Our ConfQA training is tremendously different in two ways: first, we use the dampener prompt, which reduces hallucination further by 5-11% in our empirical study; second, we focus on simple factual questions from the DBPedia knowledge graph, which increases factuality by up to 30%. Our experiments also show that requiring consistency in addition can cause a large correctness regression. We present the experimental comparison in detail in Section 5.
3 Methodology
3.1 Problem Definition
@@
The RAG Triggering problem takes an input question $Q$ and outputs a boolean regarding whether to trigger RAG: $T:Q\to\{0,1\}$. In other words,
$A=\begin{cases}RAG(Q,{\cal R}),&\text{if }T(Q)=1,\\[2.0pt] M(Q),&\text{if }T(Q)=0.\end{cases}$ (1)
An ideal triggering strategy should invoke RAG only when the LLM’s inherit answer $M(Q)$ is incorrect, thereby minimizing unnecessary overhead while optimizing quality. This reduces the probelm to train the LLM $M$ to recognize when its own output may be unreliable and to return an "Unsure" response in such cases. This response then serves as a signal to trigger RAG. We next describe ConfQA, a fine-tuning approach for calibrating model confidence, and describe how it supports our triggering strategy.
3.2 ConfQA: When to Say Unsure?
The goal of ConfQA is to fine tune an LLM to only answer a question that it has high confidence about, and admitting “I am unsure” otherwise. We have three key intuitions for this fine-tuning. First, we calibrate the LLM’s confidence by showing the ground truth. Second, we introduce a dampener prompt “Answer only if you are confident”, to explicitly guide LLM’s behavior. Third, as we wish to regularize behavior only for factual statements, we focus the teaching on atomic facts (attributes of entities) to avoid distractions of other factors.
We prepare the training data as a collection of question–label pairs, where each label provides the model with the appropriate response to generate. The questions ask for atomic facts, and are generated from DBPedia, which covers a diverse set of domains (Intuition #3). We used the open-sourced script from Sun et al. (2023a) to generate question-answer pairs from DBPedia, evenly distributed among different entity popularity: head, torso, and tail entities.
We generate labels as follows. First, we prompt Llama-3.1-70B model to answer the DBPedia-based questions (Prompt 1 in Appendix 9). Then, we prompt Llama-3.1-405B to judge if the answer is consistent with the ground truth (Prompt 2 in Appendix 9). If the answer is correct, the label is the ground truth answer; otherwise, the label is “I am unsure about the answer” (Intuition #1).
We provide the dampener prompt as the system prompt both in training and in inference, as an explicit instruction for the model to suppress hallucinations (Prompt 1 in Appendix 9). We call our fine-tuned model ConfQA, denoted by $\hat{M}$.
3.3 ConfRAG: When to Trigger RAG?
Since ConfQA is fine-tuned to answer a question only if it is confident, we can invoke the RAG pipeline when it says unsure, and rely on the LLM-generated answer otherwise.
$A=\begin{cases}RAG(Q,{\cal R}),&\text{if }\hat{M}(Q)=\text{"unsure"},\\[2.0pt] \hat{M}(Q),&\text{otherwise}.\end{cases}$ (2)
Figure 2 depicts The ConfRAG invocation architecture. For dynamic questions—those requiring up-to-date information—the system always responds through the RAG pipeline. For static questions, the system runs LLM generation ($\hat{M}$) and the RAG pipeline in parallel. If the LLM produces a valid answer, the RAG process is early-stopped and the LLM output is returned; otherwise, the system waits for and outputs the RAG result.
@@ 4.1 Benchmarks and Metrics
We experimented with three Short-form factuality benchmarks, where the answers are mostly short; the question include both simple questions regarding an attribute of an entity, and complex ones that require comparison, aggregation, reasoning, and post-processing. Table 1 summarizes the benchmarks and Appendix 10 gives details.
@@
For model metrics, following the CRAG benchmark (Yang et al., 2024), we compute the percentage of correct, incorrect (i.e., hallucinations), and missing (not attempted) answers, and take Factuality = correct% - incorrect% as our major metrics. Factuality ranges from -1 to 1 and penalizes hallucinations more than missing answers. We use prompt based LLM-as-a-judge to evaluate model answers. As observed in Yang et al. (2024), LLM-as-a-judge achieves 99% accuracy.
For triggering, we compute the precision and recall of the triggering decisions compared to the oracle solution that triggers when an answer is incorrect or missing. Take the DBPedia results in Table 2 as an example. The ground truth triggering is 48.0 (sum of Miss and Incor of Llama-3.1).
[…]
@@
We conduct experiments using six well-known LLMs: Llama3.1-8B, Llama3.1-70B (Touvron et al., 2023), GPT-4o-mini and GPT-4o (OpenAI et al., 2024), Claude3.5-Sonnet [^1] https://www.anthropic.com/news/claude-3-5-sonnet and Claude3.5-Haiku [^2] https://www.anthropic.com/claude/haiku.
Our fine-tuning uses Llama-3.1-70B as the backbone, and have observed similar results when fine-tuning Llama-3.1-8B. We conducted a simple scaling-law study (see Appendix 14) and decided to run one epoch on 3K samples of training data, with a learning rate of 1e-6, a batch size of 1, a gradient accumulation step 1. In addition to Llama models, we also fine tuned QWen2.5-7B-Instruct and Gemma-3-4B-IT. Experiment results are discussed in Appendix 17.
We conducted experiments on Nvidia H100 96GB HBM2e GPUs with different configurations. For Llama3.1-70B models, we fine tuned on 32 GPUs and inference on 8 GPUs.
4.2 Analysis of LLM’s Self-Reported Confidence (RQ1)
Figure 3: LLMs’ self-reported confidences is correlated with QA accuracy, but often over-confident.
With the three short-form benchmarks, we answer the first research questions:
  • RQ1. Does an LLM know what it knows?
We prompted the LLM to directly provide a confidence score between 0 and 1 along with its answer (prompt template in Appendix 3). We removed missing answers, divide the reported confidences by equal-sized quantile bins, and plot the average accuracy within each bin. We investigated the calibration between confidence and QA accuracy; that is, when the model has a confidence of 0.8, is the QA accuracy close to 80%? Figure 3 shows the calibration, leading to three observations. (We in addition compare calibration for head, torso, tail entities in Appendix 11 and with consistency of answers in Appendix 12).
@@ 4.2 Analysis of LLM’s Self-Reported Confidence (RQ1)
As this analysis shows, self-reported confidence tends to be over-confident so cannot serve the purpose of RAG triggering decision, justifying the need for fine-tuning.
5 Experimental Results
In this section, we study the effectiveness of ConfQA and ConfRAG, and answer the following two research questions.
  • RQ2. Can we teach LLMs to refrain from hallucinations?
  • RQ3. What is the optimal strategy for RAG triggering?
5.1 Alternative Solutions